OpenStudy (anonymous):

A physical education class has three times as many girls as boys. During a class basketball game, the girls average 18 points each, and the class as a whole averages 17 points per person. How many points does each boy score on average?

OpenStudy (ajprincess):

Let the number of boys be x. Then the number of girls will be 3x. The total points earned by girls=3x*18 =54x The total points earned by the class=17(x+3x) =17*4x =68x The total points earned by boys=68x-54x =14x Average points of boys=14x/x =14 Is that clear?

OpenStudy (ajprincess):

@ronald_m09

OpenStudy (anonymous):

how do you get the 4?

OpenStudy (ajprincess):

vch 4?

OpenStudy (anonymous):

the 4x in the equation "17(x+3x)= 17*4x"

OpenStudy (ajprincess):

Total number of children in the class= number of boys+number of girls. =x+3x =4x Is that k?

OpenStudy (anonymous):

so both of the X's added together = 1? or is the x treated as 1?

OpenStudy (ajprincess):

bt y one?

OpenStudy (ajprincess):

Let the number of boys be x. Then the number of girls will be 3x. Is this clear?

OpenStudy (anonymous):

I suppose, but you didn't answer my question

OpenStudy (ajprincess):

i didnt treat x as 1.

OpenStudy (ajprincess):

Do u knw like terms?

OpenStudy (anonymous):

not really

OpenStudy (ajprincess):

Oh k then. Like terms are terms containing the same variable with same exponents. that is x and 2x are like terms. You can add x and 2x as adding two numbers. In the same way x^2 and 3x^2 are like terms. U can do addition or subtraction with them. Bt u cant add x and x^2 or 2x and 3x^2. Is that clear?

OpenStudy (anonymous):

yeah that helps

OpenStudy (ajprincess):

according to that when u add x and 3x u get 4x. Is that clear?

OpenStudy (anonymous):

yes i understand alot better now, thanks

OpenStudy (ajprincess):

yw.:). if it is still confusiv ask me. I will try my best to clearify it.

OpenStudy (anonymous):

ok youve been a big help thank you.

OpenStudy (ajprincess):

welcome.:)